# R3R: Decentralized Multi-Agent Collision Avoidance with Infinite-Horizon Safety

### v1

@@ IV Simulation Results
As a case study, we consider 2D curvature-constrained agents operating under the Dubins [20] vehicle dynamics
@@
where $v$ is a fixed velocity, and $\omega$ is the angular acceleration input, which is bounded by $|\omega|\leq\omega_{\text{max}}$. This system is representative of a fixed-winged aircraft operating at a constant velocity. We perform simulations of up to $128$ agents, operating in different environments, with static obstacles represented by occupancy grids. All simulations were performed in julia, using Agents.jl [21], running on a 2022 MacBook Air (Apple M2, 16 GB). All simulation code, as well as scripts to recreate all figures in this paper, is made public.
The nominal planner is a Dynamic-RRT* [^4] Implementation: https://github.com/MarshallVielmetti/DynamicRRT.jl, a dubins-primitive based RRT* planner that plans dynamically feasible trajectories, while avoiding static obstacles and other agents. It is important to note that $R3R$ and gatekeeper are planner-agnostic–safety is a result of R3R and gatekeeper, not the planner.
#Agents - Environment Safety % Success %
@@
Table I: Simulation results of R3R & gatekeeper approach. Demonstrates safety and deadlock resilience.
Committed trajectories are constructed by maximizing the time for which the agent’s nominal can be tracked before switching to the backup controller. The backup controller for the Dubins dynamics was implemented as a time-parameterized loiter circle.
@@
Results from the simulations are shown in table˜I. Each scenario was run $5$ times, with all simulation parameters held constant. For the city-like and willow garage scenarios, agent starting positions and goal positions were randomized at each trial. Success was defined as the percent of agents making it to their goal positions. Safety is defined as all environmental constraints being satisfied, and no agents coming within inter-agent collision avoidance distance. The simulation parameters for the city-like environment (approx. $100\times 100$) were: $R^{\text{comm}}=16.0,\delta=0.5,R^{\text{plan}}=5.16$. For the $32$ agent simulation, agents averaged $4.7$ neighbors at each replanning iteration, with a maximum of $10$. Replanning steps took an average of 1.41.4ms.
We see R3R achieves consistently high success and safety rates across all tested scenarios, demonstrating its effectiveness in complex multi-agent environments. See fig.˜3 for an example of $64$ agents operating in the city-like environment. The only failure case was in the $128$-agent city-like scenario, where in one iteration a group of agents were unable to reach their goals within the simulation time limit. This was due to a combination of high density and limited free space, which led to deadlock situations that the current implementation could not resolve. Future work will focus on enhancing deadlock resolution strategies to further improve success rates in such challenging scenarios.
IV-1 Baseline Comparison
We compare the performance of our solver to a baseline Nonlinear MPC in Julia using Ipopt[22] and JuMP[23]. For simplicity, we demonstrated this comparison in an empty environment. Agents are given the fully-planned trajectories of other agents within their communication radius. Inter-agent collision constraints are encoded in the optimization problem as hard constraints. If the resulting problem is infeasible, we attempt to solve a relaxed version of the problem, with high penalties for constraint violation.
Due to the effects of the communication radius, the NMPC solver can find itself in unrecoverable states where no safe solution exists. It thus has a non-zero rate of safety violations, with performance degrading significantly as communication radius decreases and density increases.
$n_{\text{agents}}$ Safety %
@@ IV-1 Baseline Comparison
Table II: % Of Trials Without Safety Violations for NMPC with different numbers of agents and communication ranges.
Our empirical results highlight that the NMPC solver performed well when the communication radius was large enough to ensure a fully connected network, and the density was low enough to ensure that the problem was not too constrained. We see these results in Table˜II.
IV-2 Runtime Analysis
Let $N$ be total agents, environment area $A$, density $\rho=N/A$, and communication radius $R^{\text{comm}}$. Assuming agents are distributed uniformly, at some replanning iteration, the avg. number of neighbors within communication range is $\lambda=\rho\pi(R^{\text{comm}})^{2}$. All online costs depend on $\lambda$ (local density), not $N$, provided $\rho$ stays bounded as $N$ grows. Thus, collision checking in nominal planning and candidate validation is $O(\lambda)$, not $O(N)$ as with centralized or fully-connected approaches. For $\lambda\ll N$, this reflects significant savings.
We validate this analysis empirically. We consider a fixed number of agents $N$ operating in a square with area $d^{2}$, where $d$ is the length of the side of the square. To evaluate the effect of density on the algorithm, we fix the number of agents $N$, and vary the dimension $d$ of the environment. To ensure the simulations are reflective of the varying density, we randomize start and goal positions for each agent, and conduct $T=20$ randomized trials for each $d$. Results are shown in fig.˜4. As expected, we see that as the problem density increases the performance of the solver decreases.
@@ V Conclusion
This paper presented R3R, a novel framework that establishes a formal, geometric link between an agent’s communication range and the range within it can plan its trajectories to ensure safety. Specifically, we introduced $R$-Bounded trajectories, which remain within a radius $R$ of their starting point. By ensuring $R$ is one-third of the communication radius, we geometrically guarantee that local safety checks imply global safety. We then leverage gatekeeper to guarantee that all agents always track valid, $R$-Bounded trajectories, ensuring forward-invariant safety.
Our work demonstrates that for a broad class of systems, scalability and formal safety guarantees are not mutually exclusive goals but can be achieved simultaneously through principled geometric constraints. The efficacy of the algorithm is demonstrated via the asynchronous deconfliction of >100 agents in obstacle environments.
Our future work will focus on addressing the method’s limitations, namely to enhance the framework’s robustness, resolve deadlocks and incorporate models of network uncertainty to provide probabilistic safety guarantees under realistic communication.

### v2

@@ IV Simulation Results and Analysis
As a case study, we consider 2D curvature-constrained agents operating under the Dubins vehicle dynamics
@@
where $v$ is a fixed velocity, and $\omega$ is the angular velocity input, which is bounded by $|\omega|\leq\omega_{\text{max}}$ [20]. This system is representative of a fixed-winged aircraft operating at a constant velocity. All simulations were performed in Julia, using Agents.jl [21], running on a 2022 MacBook Air (Apple M2, 16 GB).
The nominal planner is a Dynamic-RRT* [^3] Implementation: https://github.com/MarshallVielmetti/DynamicRRT.jl, an RRT* planner based on Dubins motion primitives that plans dynamically feasible trajectories, while avoiding static obstacles and other agents. Committed trajectories are constructed by maximizing the time for which the agent’s nominal can be tracked before switching to the backup controller. The backup controller for the Dubins dynamics was implemented as a time-parameterized loiter circle, the location of which is determined as part of constructing the candidate trajectory.
Table I: Simulation results of R3R & gatekeeper approach.
#Agents - Environment Safety Success
@@
Results from the simulations are given in table˜I. Each scenario was run $5$ times. For the city-like and willow garage scenarios, agent start and goal positions were randomized in each trial. Safety is defined as all environmental constraints being satisfied, and no agents coming within inter-agent collision avoidance distance. Success was defined as the percent of agents making it to their goal positions. The simulation parameters for the city-like environment (approximately $100\times 100$) were: $R^{\text{comm}}=16.0,\delta=0.5,R^{\text{plan}}=5.16$, and maximum turning radius of $0.25$. For the $32$-agent simulation, agents averaged $4.7$ neighbors at each replanning iteration, with a maximum of $10$. Replanning steps took an average of $1.4$ ms.
The simulation results validate our theoretical safety guarantees, and the high success rates demonstrate the algorithm’s deadlock resilience, even in dense scenarios, an example of which is given in Fig.˜3. The only failure case occurred in the $128$-agent city-like scenario when a group of agents were unable to reach their goals within the simulation time limit, due to a combination of high density and limited free space. This resulted in a deadlock situation that the current implementation could not resolve. Safety was never violated. Future work will focus on enhancing deadlock resolution strategies to further improve success rates in such challenging scenarios.
IV-A Baseline Comparison
While comparison against state-of-the-art (e.g., MADER [9], DREAM [17], and EGO-Swarm [15]) is desirable, these methods fundamentally assume fully-connected communication graphs and global information sharing, and adapting them to operate under strict, distance-based communication constraints changes their core behavior. Relative to these methods, our approach sacrifices global optimality for guaranteed constraint satisfaction under partial information.
We compare the performance of our solver to a baseline Nonlinear MPC in Julia using Ipopt [22] and JuMP [23], which we demonstrate in an empty environment. Inter-agent collision constraints with respect to neighboring agents are encoded in the optimization problem as hard constraints. If the resulting problem is infeasible, we attempt to solve a relaxed version of the problem, with high penalties for constraint violation.
Due to the effects of the communication radius, the NMPC solver finds itself in unrecoverable states where no safe solution exists, and we see the rate of safety violations increase substantially as problem density increases.
Table II: Evaluation of NMPC baseline.
$n_{\text{agents}}$ Safety
@@ IV-A Baseline Comparison
IV-B Runtime Analysis
Let $N$ be the total number of agents, $A$ be the environment area, $\rho=N/A$ be the agent density, and $R^{\text{comm}}$ be the communication radius. Assuming agents are distributed uniformly, at some replanning iteration, the average number of neighbors within communication range is $\lambda=\rho\pi(R^{\text{comm}})^{2}$. All online costs depend on $\lambda$ (local density), not $N$, provided $\rho$ stays bounded as $N$ grows. Thus, collision checking in nominal planning and candidate validation is $\mathcal{O}(\lambda)$, not $\mathcal{O}(N)$ as with centralized or fully-connected approaches. This analysis is confirmed empirically by the results given in Fig.˜4.
@@ V Conclusion
This paper presented R3R, a novel framework that establishes a formal, geometric link between an agent’s communication range and the bounds within which it can plan its trajectories. Specifically, we introduced $R$-Bounded trajectories, demonstrating that restricting an agent’s plan to a function of its communication radius and collision avoidance radius geometrically guarantees that local deconfliction implies global safety. We leverage gatekeeper to ensure that all agents asynchronously construct and track valid, $R$-Bounded trajectories, achieving forward-invariant safety without global information. Our work demonstrates that for a broad class of systems, scalability and formal safety guarantees are not mutually exclusive goals but can be achieved simultaneously through principled geometric constraints. The efficacy of the algorithm is demonstrated via the asynchronous deconfliction of >100 agents in obstacle-rich environments. Our future work will focus on addressing the method’s limitations, namely to enhance the framework’s robustness, resolve deadlocks, and incorporate models of network uncertainty to provide probabilistic safety guarantees under realistic communication conditions.